LCM of w²-64 and w²+13w+40
Answer by stanbon(75887)

LCM of w²-64 and w²+13w+40
---
Factor:
w^2-64 =(w-8)(x+8)
---
w^2+13w+40 = (w+5)(w+8)
------------------------------
The LCM must have each of the different factors in their hightest power.
---
LCM = (w+5)(w-8)(w+8)
=========================
Cheers,
Stan H.